the mean of six numbers is 41 three of the numbers are 32,31 and 42 the remaining three numbers each equal to a. what is the sum of six numbers?. find the value?
Answer by hamsanash1981@gmail.com(151)

Mean of six numbers is 41, then total sum = 41*6 = 246
Given the numbers are 32,31,42 and other three are a,a,a
then, 32+31+42+a+a+a = 246
105+ 3a = 246
3a = 141
a = 47